A cylinder that contains 7kg of CO2 with molar mass 44kg/kmol, occupying a volume of 0.025 m^3 at 1.1 bar, is compressed reversibly until the pressure is 230pa. Calculate final temperature, work done on CO2, heat supplied when the process is isothermal.


1
Expert's answer
2021-05-14T09:38:02-0400

The carbon dioxide (like air, nitrogen, oxygen, hydrogen, noble gases) can be treated as an ideal gas, so, the temperature can be found from the ideal gas law:


"T=\frac{P_1V_1}{nR}=2.08\text K."

The work done by gas is


"W=P_1V_1\ln\frac{V_2}{V_1},\\\space\\
\frac{P_1V_1}{T}=\frac{P_2V_2}{T},\\\space\\
\frac{V_2}{V_1}=\frac{P_1}{P_2},\\\space\\
W=P_1V_1\ln\frac{P_1}{P_2}=16970\text{ J}."

Heat supplied:


"Q=\Delta U+W,\\\space\\
\Delta U=0,\\\space\\
Q=W=16970\text{ J}."